On 12/31/16, flowmo inc. issued 10-year, 8% bonds with a face value of $10,000,000 for a price to yield a market rate of 6.5%. what is the amount of interest expense reported on flowmo’s income statement for 2018?
Answer #1
Year end Beginning Balance Interest expenses Coupoun payment Principal Repaid Ending Balance
a b c=b*6.5% d e=d-c f=b-e
2016 0 0 0 0 $       1,10,78,325
2017 $       1,10,78,324.53 $           7,20,091.09 $ 8,00,000.00 $           79,908.91 $ 1,09,98,415.63
2018 $       1,09,98,415.63 $           7,14,897.02 $ 8,00,000.00 $           85,102.98 $ 1,09,13,312.64